OpenStudy (anonymous):

dx/dy=xy/3 y>0 Find the particular solution y=f(x) to the given differential equation with the initial condition f(0)=4. please explain!

OpenStudy (irishboy123):

separate it out and solve is it dx/dy = xy/3 or dy/dx = xy/3

OpenStudy (anonymous):

oops its actually \[\frac{ dy }{ dx}=\frac{ xy }{ ? }\]

OpenStudy (irishboy123):

Question mark?